# Quantum-enhanced Markov Chain Monte Carlo for Combinatorial Optimization

### v1

@@ I Introduction
Whilst we navigate these challenges, heuristic quantum-classical frameworks offer a practically motivated direction for near-term quantum computing research [9, 19]. For quantum optimization, methods such as warm-starting leave much to be hopeful about, even before fault-tolerance [1, 19]. Today’s quantum devices–despite inherent noise–have achieved a scale, speed, and fidelity that enable computations beyond the reach of exact classical simulation [38]. Hence, it is now possible to benchmark the capabilities of quantum-enhanced solvers on non-trivial optimization problem instances, such as the ones provided in the Quantum Optimization Benchmarking Library (QOBLIB) [39]. This resource provides a framework consisting of combinatorial optimization problems, benchmarking metrics, and baseline results to assist the community in exploring the path to quantum advantage in optimization. This also enables us to better understand the potential practical usefulness of quantum optimization algorithms in addressing problems that map to industry-relevant applications.
As quantum hardware becomes increasingly capable, we can begin to investigate theoretical results with non-trivial problem instances. One set of interesting theoretical and early numerical results include those from the quantum-enhanced Markov chain Monte Carlo (QeMCMC) algorithm, originally developed for approximating challenging probability distributions [43, 41]. The authors in Ref. [41] showed that with a quantum proposal distribution, QeMCMC demonstrated up to a quartic advantage in reaching the target distribution more efficiently compared to classical Markov chain Monte Carlo (MCMC). Subsequent work has further investigated behavior, limitations, and variants of QeMCMC [48, 49, 26, 12, 10].
@@
We structure this work as follows. In Section II, we provide background material before presenting key methods used herein in Section III, including QeMCMC, warm-starting quantum optimization and parallel tempering. We explain how these methods work together to form the basis of our workflow in Section III.6, accompanied by Figure 1. In Section IV.2, we test the applicability of applying the QeMCMC-inspired optimization algorithm to MIS. We provide comparisons to known classical methods in Section IV.3, discuss anticipated scaling, or feasibility with increasing problem size, of our algorithm in Section IV.4 and conclude in Section V.
@@ III.3 Parallel tempering
    Local MCMC updates: Each replica $i$ performs a number of standard MCMC updates at its own temperature.
@@ III.6 Experimental Workflow
We repeat this loop for a maximum of $200,000$ iterations, or until convergence to the target objective. Our algorithm has many parameters, which we tune using a combination of analytical guidelines drawn from the literature, in addition to heuristic tuning. These parameters differ for each problem instance. A table of the relevant optimizable parameters is presented in Table 1. This tuning of parameters strongly influence mixing efficiency and convergence [16, 40, 47, 11]. Further details regarding parameter tuning for each MIS problem instance is explored further in Appendices C and D.
@@ IV Results
We study three different problem instances, most of which are taken from the QOBLIB [39], executed both in simulation and on IBM quantum hardware. These instances are derived from real-world animal social interaction networks, with qubit or node counts $n\in\{17,52,117\}$ and corresponding edge counts $m\in\{91,516,1027\}$. These graphs share a common provenance and representation as simple, undirected interaction networks. As social networks, they exhibit heterogeneous degree distributions and nontrivial interaction structure, while varying in size and density. Together, these instances provide a natural progression in problem size for probing the scalability of our algorithm. In the remainder of this Section, we focus on the $117$ node problem.
IV.1 Simulation results
The algorithm is used to solve MIS problems both using matrix product state (MPS) simulation and on IBM quantum hardware. Simulations were conducted with the Qiskit Aer MPS simulator, with a truncation thresholds between of $1\times 10^{-6}$ and $1\times 10^{-10}$. This threshold is defined as the number of Schmidt coefficients for which the sum of their squares is smaller than the threshold [36]. Outside of this threshold, values are discarded.
@@ IV.1 Simulation results
% convergence
across 10 repeats
Median
convergence
1 0.01 10.0% 863
2 0.11 30.0% 796
3 0.21 20.0% 706
4 0.51 10.0% 619
5 1.01 0%
@@
Table 2: Replica convergence statistics for 5 replicas across 10 independent repeats of simulated experiment for 117 node graph problem instance.
In Table 2, we see the number of iterations needed to converge to the target objective (within a maximum of $200,000$ iterations). These results are also visualized in Appendix D.
We see that colder replicas have a higher rate of success in converging to their target values, but that a warmer replica, Replica $4$, was able to converge in the fewest number of iterations in total, likely due to its more broader search behavior.
IV.2 Hardware results
When running on real quantum hardware, the subsampling procedure used to mitigate against noisy samples for each of these problem instances, is defined in Table. 3. In Figure 2, we see the global optimum is recovered after $151$ iterations on the hardware. $10,000$ shots are taken per iteration, followed by randomly subsampling from the best ten samples, using the CVaR method presented in Section III.4. For the winning replica, represented in orange, the Hamming distance between accepted solutions at each iteration is displayed.
Experimental results show that the quantum hardware exceeds expectations in terms of algorithm performance indicated from our classical MPS simulations, presented in Table 2, since convergence was achieved in fewer iterations on hardware than simulation. This is likely due to truncation error dominating in the noiseless MPS simulation for such a large system, since we observe worsening performance with increased truncation error introduced to the MPS experiments via controlled testing.
@@ IV.2 Hardware results
Iterations
to converge
@@
ibm_aachen 17 - $740$ $96$ 32 5 5 Best sample taken
ibm_aachen 17 - $740$ $96$ 2 55 $10,000$ Sample from best ten
ibm_aachen 52 13 $1964$ $84$ 81 $41\times 10^{9}$ $10,000$ Sample from best ten
ibm_boston 52 13 $1964$ $84$ 17 $24,503$ $10,000$ Sample from best ten
@@
Table 3: Summary of notable experimental results. All experiments operated with 55 replicas. The estimated number of shots per iteration are derived from the logic explained in Section III.4, where differences in how the devices experience noise results in different values for estimated numbers of shots required per iteration. At the time of writing, error rates were lowest for the ibm_boston device, followed by ibm_pittsburgh and ibm_aachen.
A table of all hardware results for the problem instances of 17, 52 and 117 nodes can be found in Table 3. We report the maximum number of SWAP layers allowed for the simplified ansaetz for these problems, alongside the associated 2-qubit gate counts and depths. We also provide the estimated upper limit on samples required to compensate the impact of noise, in addition to the number of samples actually taken where this estimated value is unfeasible on hardware, alongside the number of iterations to converge.
IV.3 Baseline comparison studies
@@
Figure 3: Comparison of experimental setups for 117 node problem instance experiments. Simulations of the quantum-enhanced algorithm are compared to classical MCMC methods used to find the target objective. In the top plot, we present an MPS simulation of the quantum-enhanced algorithm. Across ten repeats, the target objective is found after a median of $\sim 90,000$ iterations for the coldest replica, with 11-shot per iteration. The middle plot shows a comparison to a classical MCMC simulation, which also uses 11-shot per iteration. Across ten repeats, the target objective is found in 4,0008,0004,000-8,000 iterations for the coldest replica, with a median of $\sim 6,000$ shots, or iterations. Finally, in the bottom plot, we display another classical MCMC simulation, where across ten repeats, the target objective is found in 4,0008,0004,000-8,000 iterations for the coldest replica, with a median of $\sim 5,000$ iterations, in this case using $10,000$ shots per iteration and a random sample taken from the best ten proposals.
The most natural comparison for our heuristic QeMCMC algorithm, is to first understand the capabilities of classical replica exchange MCMC. In this Section, we compare results for the 117 node instance. There exist various possible metrics to compare these experiments, including number of samples, iterations, or wall-clock time. It is usual for standard MCMC to take one sample per iteration, and evaluate this proposal using the MH criteria in Eq. (4). We begin by comparing results in this setting. The top plot in Figure 3 demonstrates our quantum algorithm using 11-shot per iteration with a noiseless MPS simulation. We use classical simulation since the anticipated runtime using quantum hardware was excessive and considered not a sensible use of resources for the 11-shot regime, since each sample is likely not noise-free.
In these simulations, experiments converge in a median of $\sim 90,000$ shots across ten trials, where median was the most favorable metric for convergence for both classical and quantum implementations. This is reasonably efficient considering the results presented for the experiment on the ibm_boston quantum device in Figure 2 took at least $1.51\times 10^{6}$ shots (for the winning replica). The performance seen in this top plot of Figure 3 is easily outperformed in $\sim 6,000$ shots required by standard MCMC, presented in the middle plot of this same Figure 3. Since we are limited from testing this setup on hardware, alternative setups are tested which provide a valid point of comparison, in terms of iterations of the algorithm where each iteration is more costly in number of shots. This setup makes better use of scarce quantum computing resources. An example is seen in Figure 2, where we take $10,000$ shots per iteration of the algorithm.
The number of shots taken per iteration is significantly less than the estimated number of shots required to compensate for noise, reported in Table 3. However, since these estimations are upper bounds (see Section III.4), in practice, we tune the number of shots such that we can expect a reasonable anticipated runtime on device, usually to $10,000$ shots per iteration. From these $10,000$ shots, we randomly subsample from the best ten proposals, as described in Section III.4 and III.6. Simulated convergence metrics of this setup are presented in Table 2, where convergence is reported in $\sim 600-900$ iterations. We observe the results from the quantum hardware require far fewer iterations to converge, see Figure 2, where the experiment converges in $151$ iterations compared to the median convergence reported of $796$ for the same replica across ten repeats in classical simulation of the quantum experiment. Presumably, this is because truncation error in this regime becomes more detrimental than hardware noise, since it is observed that performance worsens with increased truncation threshold.
For completeness, we also allow this equivalent experimental setup with the classical proposal distribution, using classical MCMC replicas. We take the exact same setup with $10,000$ shots per iteration and apply this to the classical case, with random subsampling from the best ten ranked solutions. This result is presented in the bottom plot in Figure 3, where we see convergence is comparable in terms of number of iterations to the classical MCMC results for the 11-shot experimental setup, shown in the middle plot of Figure 3. This gives us confidence that this adapted experimental setup is reasonable. Nevertheless, results from this classical MCMC simulation shown in the middle plot of Figure 3 converge within 4,0008,0004,000-8,000 iterations, with a median of $\sim 5,000$ iterations. This is an order of magnitude larger in terms of number of iterations than the quantum-enhanced algorithm, both in simulation and with regards to results from the real quantum hardware.
Additionally, variations of 11- and 22-distance local searches were tested, which were considered standard baseline comparisons for this type of optimization algorithm. None of these converged within $200,000$ allowed iterations. Various combinations of standard MCMC tests were explored, including up to $25$ bitflips per sample for each replica, as informed by the Hamming distance between subsequent accepted solutions, $s$, on the quantum hardware, see Figure 2. The most competitive classical comparison was the classical MCMC setup described above, with up to $5$ bitflips allowed per iteration. Finally, we explored alternative quantum-enhanced variants of the algorithm, including an equivalent non-warm started $p=2$ QAOA approach, which suffered from a finite size effect beyond the 17 node problem instances. We also tested a warm-started $p=0$ setup, where circuits are composed purely of the (warm-started) initial state only, with no cost or mixer layers. Again, none of these tests were seen to converge to the global optimum.
@@
Figure 4: A summary of anticipated performance scaling in terms of shots and iterations of the classical vs. quantum-enhanced algorithm. Median iterations to converge are reported for the 17, 52 and 117 node graphs after ten repeated tests of the classical and quantum-enhanced algorithms. A line of best fit provides us with a naïve indication of scaling behavior as to how we anticipate each method to perform with increasing problem instance size. On the left, we present anticipated performance scaling in terms of shots for the classical vs. simulated quantum-enhanced algorithm, using the Qiskit Aer MPS simulator. On the right, we see anticipated performance scaling in terms of iterations of the classical vs. quantum-enhanced algorithm, both simulated and run on real hardware.
All comparison studies were initialized from the same starting point, which was drawn from sampling the non-warm started $p=2$ QAOA circuit. The studied problem instances have been solved to optimality using CPLEX, which provides reference target objectives represented across Figures documented in this paper [35].
IV.4 Scaling analysis
In the previous Section IV.3, we discuss the difficulty in comparing classical and quantum-enhanced MCMC approaches within this parallel tempering framework for solving optimization problems. Classical MCMC approaches perform well in a 11-shot per iteration experimental setup, whereas the quantum-enhanced case performs more efficiently if we take many shots per iteration. Therefore, it is useful to look instead at the scaling behavior of both approaches, looking at both number of samples and number of iterations, as we aim to solve larger and more complex problem instances.
To understand whether there is interesting systematic behavior for solving optimization problems of increasing size with both the classical and quantum-enhanced MCMC approaches in this parallel tempering framework, we refer to Figure 4. Although only three related problem instances have been studied, this plot provides a very early indication of how we might anticipate performance scaling for the algorithm presented in this manuscript, with increasing problem instance size as a proxy metric for problem difficulty.
For the classical and simulated quantum-enhanced algorithm, ten independent trials were run and various statistics computed for each of the three problem instances represented in the Figure 4. The median values are plotted for both number of shots and/or iterations as an indication of convergence rate for these implementations. In the first subplot of Figure 4, with the 11-shot experimental setup, we see the classical MCMC algorithm outperforming the simulated quantum-enhanced experiment for each of the three problem instances. As mentioned previously, it was not possible to run this experiment on hardware as the anticipated time required on the quantum device for this 11-shot setup was unfeasible. However, we are primarily interested in the slopes of the fitted lines in Figure 4, which translate to the exponential growth rates. In the left-hand plot of Figure 4, the gradient of the line for the MPS simulation of the quantum algorithm grows at $\sim 86\%$ of the classical rate, suggesting early evidence of a scaling advantage of the quantum-enhanced algorithm, at least for the three studied problem instances. In the right-hand subplot of Figure 4, $10,000$ shots are taken from each iteration. We randomly subsample from the best solutions to compensate for noise, as described in Section III.4. We see that the fitted line for the classical non-standard MCMC implementation is more performant only with the smaller problem sizes.
Three generations of the IBM quantum Heron devices were tested for each of the problem instances here, specifically Heron r1 (ibm_torino), Heron r2(ibm_aachen) and Heron r3 (ibm_boston), where Heron r3 devices have the lowest error rates of the IBM Quantum superconducting fleet at the time of writing. The relevant data for each device is available via the IBM Quantum platform.
@@ IV.4 Scaling analysis
Generally, the gradients of the fitted lines in the right-hand side plot of Figure 4 are worse than on the left plot. Yet, the gradient of the red line, on the right of Figure 4, representing the Heron r3 device results, has a slightly improved gradient to those reported on the left. These Heron r3 hardware results outperform both the equivalent classical MCMC test, and the MPS simulated quantum algorithm. Again, this suggests a promising scaling property for the quantum algorithm with larger problem instances, both in comparison to its predecessor hardware and with regards to both the classical and simulated quantum approaches. Interestingly, for the $52$ node problem on the right of Figure 4, the MPS simulation performs best. However, for the $117$ node instance, the MPS simulation is outperformed by the real Heron r3 hardware. At the same time, with reduced hardware noise, the quantum device outperforms the simulation. From the data provided so far, we see evidence of promising scaling behavior of our quantum-enhanced algorithm with respect to problem size. With larger and improved quantum hardware, we intend to collect data for larger problem instances to understand which approach continues to show better scaling behavior, since we acknowledge only three non-trivial problem instances were studied in this manuscript. More data is required to draw stronger conclusions.

### v2

@@ I Introduction
Whilst we navigate these challenges, heuristic quantum-classical frameworks offer a practically motivated direction for near-term quantum computing research [9, 19]. For quantum optimization, hybrid quantum-classical algorithms and techniques such as warm-starting leave much to be hopeful about, even before fault-tolerance [53, 1, 19]. Today’s quantum devices–despite inherent noise–have achieved a scale, speed, and fidelity that enable computations beyond the reach of exact classical simulation [39]. Hence, we can now benchmark the capabilities of quantum-enhanced solvers on non-trivial optimization problem instances, such as the ones provided in the Quantum Optimization Benchmarking Library (QOBLIB) [40]. This resource provides combinatorial optimization problem instances, benchmarking metrics, baseline results to assist the community in exploring the path to quantum advantage in optimization, as well as a foundation to better understand the potential practical usefulness of quantum optimization algorithms in addressing industry-relevant applications.
As quantum hardware becomes increasingly capable, we can begin to investigate theoretical results with non-trivial problem instances. One set of interesting theoretical and early numerical results include those from the quantum-enhanced Markov chain Monte Carlo (QeMCMC) algorithm, originally developed for approximating challenging probability distributions [44, 42]. The authors in Ref. [42] showed that with a quantum proposal distribution, QeMCMC demonstrated up to a quartic advantage in reaching the target distribution more efficiently compared to classical Markov chain Monte Carlo (MCMC). Subsequent work has further investigated behavior, limitations, and variants of QeMCMC [49, 50, 26, 12, 10, 34].
@@
We structure this work as follows. In Section II, we provide background material before presenting key methods used herein in Section III, including QeMCMC, warm-starting quantum optimization and parallel tempering. We explain how these methods work together to form the basis of our workflow in Section III.6, accompanied by Figure 1. In Section IV.3, we test the applicability of applying the QeMCMC-inspired optimization algorithm to MIS. We provide comparisons to known classical methods in Section LABEL:sec:comparison, discuss anticipated scaling, or feasibility with increasing problem size, of our algorithm in Section LABEL:sec:scaling and conclude in Section V.
@@ III.3 Parallel tempering
    Local MCMC updates: Each replica $i$ perform standard MCMC updates at its own temperature.
@@ III.6 Experimental Workflow
We repeat for a maximum of $200,000$ iterations, or until convergence. Tuning algorithm parameters for each problem instance strongly influences mixing efficiency and convergence [16, 41, 48, 11]. We use analytical guidelines drawn from the literature, as well as heuristic tuning. A summary of these parameters is presented in Table 1, and further details regarding their optimization can be found in Appendices C and D.
@@ IV Results
We study five different problem instances sourced from the Network Repository, a large public collection of graph datasets used in machine learning, network science, and graph algorithm benchmarking [51]. These graphs are derived from real-world animal social interaction networks, with qubit or node counts $n\in\{17,39,52,80,117\}$ and corresponding edge counts $m\in\{91,245,454,714,1028\}$. The graphs share a common provenance and representation as simple, undirected interaction networks. As social networks, they exhibit heterogeneous degree distributions and nontrivial interaction structure, while varying in size and density. Together, these instances provide a natural progression in problem size for probing the scalability of our algorithm. Many of these problem instances are also included in the QOBLIB [40].
We focus our analysis primarily on the $117$-node problem instance, which consists of a fully connected component of $110$ nodes and $7$ isolated nodes. While the isolated nodes could be removed, we keep the full graph without any pre-processing. We explore these problem instances using classical replica-exchange MCMC, classical matrix product state (MPS) simulation of the QeMCMC algorithm, and executions on real IBM quantum hardware. We first present classical baseline results using a classical proposal distribution, which then serve as a benchmark for our quantum-enhanced approach. Then, we run our algorithm using both classical MPS simulation and real quantum hardware. Experimental details such as number of replicas, tuning of the temperature ladder (described in Appendix D) and regularity of replica exchanges are kept consistent across setups. For classical MCMC and MPS simulations, we compare both single-shot and multi-shot settings per iteration. In contrast, for real quantum hardware, we only consider multiple shots per iteration to account for noise, as motivated in Section III.4. The globally optimal objective values for all problem instances were computed via CPLEX, an advanced classical optimization suite capable of solving the described problem instances to optimality [36].
IV.1 Classical baseline results
The most natural comparison for our heuristic QeMCMC algorithm, is to first understand the capabilities of classical replica exchange MCMC. In this setting, the proposal distribution is constructed via classical bitflips. Various proposal strategies for the classical MCMC setup were explored in this work, including uniformly sampled bitflip proposals of up to $25$ bitflips per sample for each replica. The most competitive classical configuration was uniform sampling with up to $5$ bitflips per iteration, for which results are presented here. We test the classical setup with both $1$ shot and $10,000$ shots per iteration. In the case where $10,000$ shots are taken per iteration, we randomly subsample from the best ranked ten proposals.
Just as is done for the quantum-enhanced algorithm setting, samples for the first iteration for each experiment are taken from the non-warm started $p=2$ QAOA circuit. This experiment is repeated $10$ times, with $200,000$ iterations per run.
@@ IV.1 Classical baseline results
Median num. iterations
until convergence
(1 shot per iteration)
Median num. iterations
until convergence
($10k$ shots per iteration)
1 0.01 6293 5101
2 0.11 7554 5051
3 0.21 8191 6447
4 0.51 8576 9196
5 1.01
@@
Table 2: Replica convergence statistics for 5 replicas across 10 independent repeats of the classical replica exchange MCMC experiment for 117 node graph instance, where either 11 shot or $10,000$ shots are taken per iteration of the algorithm, with random subsampling from the best ten proposals.
The lowest number of samples required to converge to the global optimum in the $1$-shot per iteration setting is $6,293$ samples or iterations. With $10,000$ shots per iteration, the lowest number of iterations required to converge to the global optimum is $5,101$ iterations. The overall results for both settings and all temperatures are shown in Table 2. While these numbers are close in terms of iterations, the total number of shots is significantly larger in the multi-shot setting. Thus, the additional shots seem to create little additional value in this noise-free setting.
IV.2 Simulation of quantum hardware
Simulations of the quantum-enhanced algorithm, with quantum circuits composing the proposal distribution, were performed using the Qiskit Aer MPS simulator. MPS simulations are subjected to a truncation threshold, which specifies the tolerance for discarding Schmidt coefficients: coefficients are truncated when the sum of the squares of the discarded values falls below the specified threshold [37]. Truncation thresholds ranged from $10^{-6}$ to $10^{-10}$ across all problem instances studied, with a threshold of $10^{-6}$ applied to simulations of the $117$ node problem, for which results are presented below. This threshold was chosen to balance minimizing truncation error against feasible runtimes for the simulations. The quantum circuits implemented via MPS are the same ansatz circuits constructed for the hardware experiments, for which the process is described in Appendix E. Again, we test both 1-shot setting and 10,000-shot setting with subsampling from the best ten proposals.
Circuits are initialized in the same way, with initial samples drawn from the non-warm started $p=2$ QAOA circuit. These experiments are repeated $10$ times, with up to $200,000$ iterations.
Replica Temperature
Median num. iterations
until convergence
(1 shot per iteration)
Median num. iterations
until convergence
($10k$ shots per iteration)
1 0.01 91668 863
2 0.11 796
3 0.21 706
4 0.51 619
5 1.01
Table 3: Replica convergence statistics for 5 replicas across 10 independent repeats of the simulated quantum experiment for 117 node graph instance, where either 11 shot or $10,000$ shots are taken per iteration of the algorithm, with random subsampling from the best ten proposals.
In Table 3, we see the number of iterations needed to converge to the global optimum for both $1$ shot per iteration and $10,000$ shots per iteration. For $1$ shot per iteration, only the coldest replica managed to converge to the optimal solution after $91,668$ samples or iterations. For $10,000$ shots per iteration, we see that colder replicas have a higher rate of success in converging to their target values, but that a warmer replica, Replica $4$, was able to converge in the fewest number of iterations in total, likely due to its broader search behavior. These results are also visualized in Appendix D. It is interesting that taking multiple samples per iteration with subsampling now strongly affects the number of required iterations to converge, unlike in Section IV.1. We expect this stems from the fact that QAOA-based proposal distributions, while introducing a bias towards good solutions, also produce a tail of much worse ones. Subsampling effectively isolates the high-quality tail of the proposal distribution, making it beneficial beyond merely mitigating noise.
IV.3 Quantum hardware results
Finally, we present the results of our heuristic QeMCMC optimization algorithm using IBM quantum hardware. In this setting, we do not explore the $1$-shot regime due to the presence of noise and because it would lead to prohibitively long runtimes. Including circuit preparation overheads on the quantum hardware, taking $1$ shot per circuit takes almost as long as taking $10,000$ shots per circuit. Thus, we execute experiments taking $10,000$ shots per iteration followed by the subsampling procedure described in Sec. III.4.
@@ IV.3 Quantum hardware results
Median num. iterations
until convergence
@@
ibm_aachen 17 $14\text{*}$ $740$ $96$ 32 5 5 Best sample taken
ibm_aachen 17 $14\text{*}$ $740$ $96$ 2 55 $10,000$ Sample from best ten
ibm_boston 39 32 $3616$ $198$ 12 $2,065$ $10,000$ Sample from best ten
ibm_boston 52 13 $1964$ $84$ 22 $24,503$ $10,000$ Sample from best ten
@@
Table 4: Summary of notable experimental results. All experiments used 55 replicas. Three repeats were run for each experiment on the hardware devices. Estimated number of shots per iteration are derived from the logic explained in Section III.4, where differences in how devices experience noise lead to different values for estimated numbers of shots required per iteration. At the time of writing, error rates were lowest for the Heron r1 ibm_boston device, followed by Heron r1 ibm_pittsburgh and Heron r2 ibm_aachen.
* This is the maximum number of SWAP layers for this graph i.e. the full graph is implemented for this instance.
Results for the different problem instances are shown in Table 4. These include the maximum number of SWAP layers allowed for the simplified ansaetz for these problems, alongside the associated 2-qubit gate counts and depths of the circuits. We report the total shots taken alongside the number of samples kept for subsampling for different instances and quantum devices, together with their convergence analysis. In both sets of experiments using the $117$ node graph, $10,000$ shots were taken on the device, with the experiments converging to the global optimum on average in $151$ and $83$ iterations respectively. The number of shots taken is significantly less than the estimated number required to compensate for noise. However, since these estimations are upper bounds (see Section III.4), in practice, we choose $10,000$ shots such that we can expect a reasonable runtime on the device.
Results for the 117-node problem instance on ibm_boston are visualized in Figure 2. The winning replica that found the optimal solution first is represented in orange. For this replica, the Hamming distance between accepted solutions at each iteration is also displayed. From the data shown in Table 4, we see that the quantum hardware exceeds expectations in terms of algorithm performance indicated from our classical MPS simulations reported in Table 3, since convergence was achieved in $151$ iterations, which is far fewer iterations on hardware than $796$ iterations predicted for the same replica via simulation. Presumably, this is because truncation error in this regime becomes more detrimental than hardware noise, since we also observed worsening performance with increased truncation error introduced to the MPS experiments via controlled testing.
IV.4 Comparison and scaling analysis
@@
Figure 3: Comparison of experimental setups for 117 node problem instance experiments. Simulations of the quantum-enhanced algorithm are compared to classical MCMC methods used to find the target objective. In the top panel, we present an MPS simulation of the quantum-enhanced algorithm. Across ten repeats, the target objective is found after a median of $\sim 90,000$ iterations for the coldest replica, with $1$ shot per iteration. The middle panel shows a comparison to a classical MCMC simulation, which also uses 11 shot per iteration. Across ten repeats, the target objective is found in 4,0008,0004,000-8,000 iterations for the coldest replica, with a median of $\sim 6,000$ shots, or iterations. Finally, in the bottom panel, we display another classical MCMC simulation, where across ten repeats, the target objective is found in 4,0008,0004,000-8,000 iterations for the coldest replica, with a median of $\sim 5,000$ iterations, in this case using $10,000$ shots per iteration and a random sample taken from the best ten proposals.
In this Section, we compare results of the different algorithms. There exist various possible metrics to compare these experiments, including total number of samples, iterations, or wall-clock time, as well as the corresponding scaling behavior across problem sizes.
First, we compare the performance of the algorithms for the 117-node problem. In Figure 3, the top panel demonstrates our quantum algorithm using a MPS simulation with $1$ shot per iteration. The experiment converges in a median of $\sim 90,000$ samples (or iterations) across ten trials, as can be seen in Table 3. This is easily outperformed in $\sim 6,000$ samples (or iterations) required by the standard $1$-shot MCMC shown in the middle panel of Figure 3 and summarized in Table 2. Finally, we take $10,000$ samples per iteration and subsample from the best ten. In the bottom panel of Figure 3, we see the classical MCMC implementation of this setting. Although convergence is comparable to the standard $1$-shot setup, the best reported result of $\sim 5,000$ iterations to converge is an order of magnitude larger than the quantum results, both simulated (see Table 3) and on device (see Table 4). We see the classical approach performs well with $1$-shot per iteration, whereas the quantum algorithm performs better if we take many shots. To understand how the algorithms scale in terms of both iterations and samples, we probe the performance of five problem instances.
@@
Figure 4: A summary of anticipated performance scaling in terms of shots and iterations of the classical vs. quantum-enhanced algorithm. Median iterations to converge are reported for the 17, 39, 52, 80 and 117 node graphs after ten repeats of the classical and quantum-enhanced algorithms, and three repeats on hardware. A line of best fit provides us with a naïve indication of scaling behavior as to how we anticipate each method to perform with increasing problem instance size. On the left, we present anticipated performance scaling in terms of shots for the classical vs. simulated quantum-enhanced algorithm, using the Qiskit Aer MPS simulator. On the right, we see anticipated performance scaling in terms of iterations of the classical vs. quantum-enhanced algorithm, both simulated and run on real Heron r1, Heron r2 and Heron r3 IBM quantum hardware.
For the classical and simulated quantum algorithms, ten independent trials were run and various statistics were computed for each of the problem instances, as shown Figure 4. Median values, equations of the fitted lines, and the quality of these fits are plotted for both number of samples and/or iterations as an indication of exponential growth rates for the five problem instances studied. On the left of Figure 4, in the $1$-shot setting, we observe the classical algorithm outperforms the simulated quantum experiment for each problem instance. However, the fitted line representing the exponential growth rate for the MPS simulations is $\sim 83\%$ of the classical results, suggesting an early indication of a scaling advantage of the quantum algorithm for the studied problems.
In the right panel of Figure 4, $10,000$ shots are taken for each iteration of the various experiments. In addition to the classical MCMC experiments with multiple samples and the MPS simulation of the quantum algorithm, which are each repeated ten times, we also run three repeats of the quantum experiments on three generations of the IBM quantum Heron devices for each of the problem instances. Specifically, we use Heron r1 (ibm_torino), Heron r2 (ibm_fez) and Heron r3 (ibm_boston), which has the lowest error rates of these devices. The relevant error data is available via the IBM Quantum platform.
@@
Most slopes of the fitted lines in the right panel of Figure 4 are worse than those on the left panel. Yet, the gradient of the red line on the right panel of Figure 4, representing the Heron r3 results from ibm_boston, has a comparable gradient to the fitted lines on the left panel. These Heron r3 hardware results outperform both the equivalent classical MCMC test, and the simulated quantum results in this multi-shot setting. Again, this suggests a promising scaling property for the quantum algorithm for the largest problem instances, both in comparison to its predecessor hardware, and with regards to both the classical and simulations of the quantum approach. Interestingly, MPS simulations perform best for the $52$ node problem on the right-hand panel of Figure 4. However, for the $80$ and $117$ node instances, the MPS simulations are outperformed by the real Heron r3 hardware. These results provide initial evidence that our quantum algorithms scales favorably with problem size compared to similar classical algorithms. Additional experiments are needed to further support this conclusion.